OpenStudy (anonymous):

Find all differentiable functions defined for x>0 which satisfy ƒ(xy)=ƒ(x)+ƒ(y), x,y>0.

OpenStudy (anonymous):

logs work, but it says "find all" so we need a proof of some kind

OpenStudy (anonymous):

just a couple quick thoughts. if \[f(xy)=f(x)+f(y)\] then \[f(x^n)=nf(x)\] for sure, and also \[f(1)=0\] (from the previous consideration)

OpenStudy (anonymous):

Is there a name/type of function that we could use to prove this?

OpenStudy (anonymous):

Any idea, PsychoTink?

OpenStudy (anonymous):

lol